Agassi Company uses a job order cost system in each of its three manufacturing departments. Manufacturing overhead is applied to jobs on the basis of direct labor cost in Department D, direct labor hours in Department E, and machine hours in Department K.

In establishing the predetermined overhead rates for 2020, the following estimates were made for the year.

Department

D

E

K

Manufacturing overhead $770,000 $1,500,000 $780,000
Direct labor costs $1,100,000 $1,875,000 $675,000
Direct labor hours 150,000 125,000 60,000
Machine hours 600,000 750,000 120,000


During January, the job cost sheets showed the following costs and production data.

Department

D

E

K

Direct materials used $210,000 $189,000 $117,000
Direct labor costs $180,000 $165,000 $56,250
Manufacturing overhead incurred $148,500 $186,000 $118,500
Direct labor hours 12,000 16,500 5,250
Machine hours 51,000 67,500 10,480

(a)

Compute the predetermined overhead rate for each department. (Round answers to 2 decimal places, e.g. 12.50 or 12.50%.)

Overhead rate
Department D %
Department E

$

per direct labor hour
Department K

$

per machine hour
Answer #1
Statement showing Computations
Particulars Amount
Estimated Manufacturing overhead            770,000.00
Estimated Direct labor costs        1,100,000.00
predetermined overhead rate of Department D=770,000/1100,000 70.00%
Particulars Amount
Estimated Manufacturing overhead        1,500,000.00
Estimated Direct labor hours            125,000.00
predetermined overhead rate of Department E = 1500,000/125,000                      12.00
Particulars Amount
Estimated Manufacturing overhead            780,000.00
Estimated machine hours            120,000.00
predetermined overhead rate of Department K = 780,000/120,000                         6.50